# TUTORIAL ARTICLE Contraction Theory for Nonlinear Stability Analysis and Learning-based Control: A Tutorial Overview

### v7

@@ 7.3 Tube-based State Constraint Satisfaction
7.4 Safe Exploration and Learning of Disturbances
All the theorems presented so far assume that the sizes of unknown external disturbances are fixed (i.e., $\bar{d}$ in Theorem 2.4 and $(\bar{g}_{0},\bar{g}_{1})$ in Theorem 2.5 are known and fixed). Since such an assumption could yield conservative state constraints, e.g., in utilizing motion planning of Lemma 7.1 with the tube of Theorem 7.1, we could consider better estimating the unknown parts, $d(x,t)$ in Theorem 2.4 or $G_{0}(x,t)$ and $G_{1}(x,t)$ in Theorem 2.5, also using contraction theory and machine learning. For example, it is demonstrated in Nakka et al. [2021] that the stochastic bounds of Theorem 2.5 can be used to ensure safe exploration and reduction of uncertainty over epochs in learning unknown, control non-affine residual dynamics. Here, its optimal motion plans are designed by solving chance-constrained trajectory optimization and executed using a feedback controller such as Theorem 4.2 with a control barrier function-based safety filter Ames et al. [2014], Kolathaya and Ames [2019]. It is also shown in Lakshmanan et al. [2020], Gahlawat et al. [2021], Sun et al. [2021] that the state-dependent disturbance $d(x)$ in Theorem 2.5 can be learned adaptively with the Gaussian process while ensuring safety.
These techniques are based on the robustness and stability properties of contraction theory-based control and estimation schemes, which guarantee their state trajectories to stay in a tube around the target trajectory as in Theorems 2.42.55.2, and 5.3, and at the same time, enable safely sampling training data for learning the unknown part of the dynamics. Sections 8 and 9 are for giving an overview of these concepts in the context of contraction theory-based adaptive control Tsukamoto et al. [2021], Lopez and Slotine [2021, 2021] and robust control synthesis for learned models Shi et al. [2019], Singh et al. [2020], Boffi et al. [2020], Manchester et al. [2021], Blocher et al. [2017].
@@ 7.4 Safe Exploration and Learning of Disturbances
8 Learning-based Adaptive Control
@@
where $x$, $u$, $x_{d}$, and $u_{d}$ are as defined in (6.1) and (6.3), and ${f}:\mathbb{R}^{n}\times\mathbb{R}^{c}\mapsto\mathbb{R}^{n}$ and $B:\mathbb{R}^{n}\times\mathbb{R}^{c}\mapsto\mathbb{R}^{n\times m}$ are known smooth functions with the uncertain parameter $\theta\in\mathbb{R}^{c}$. Due to Theorems 2.4 and 5.2, we can see that the robust control techniques presented earlier in Theorems 4.24.66.4, and 6.5 are still useful in this case if the modeling errors $\|f(x,\theta)-f(x,\theta_{n})\|$ and $\|B(x,\theta)-B(x,\theta_{n})\|$ are bounded, where $\theta_{n}$ is a nominal guess of the true parameter $\theta$. However, there are situations where such assumptions are not necessarily true.
We present a method of deep learning-based adaptive control for nonlinear systems with parametric uncertainty, thereby further improving the real-time performance of robust control in Theorems 5.2 and 5.3 for model-based systems, and Theorems 9.1 and 9.2 for model-free systems. Although we consider continuous-time dynamical systems in this section, discrete-time changes could be incorporated in this framework using Boffi et al. [2020], El Rifai et al. [2005]. Also, note that the techniques in this section Tsukamoto et al. [2021] can be used with differential state feedback frameworks Manchester and Slotine [2017], Singh et al. [2017, 2020], Wang et al. [2019, 2020], Manchester and Slotine [2018] as described in Theorem 4.6, trading off added computational cost for generality (see Table 3 and Lopez and Slotine [2021, 2021]).
8.1 Adaptive Control with CV-STEM and NCM
@@
leading to the following theorem Tsukamoto et al. [2021] for the NCM-based adaptive control. Note that we could also use the SDC formulation with respect to a fixed point as delineated in Theorem 3.2 Tsukamoto and Chung [2021], Dani et al. [2015], Tsukamoto and Chung [2019].
@@
Suppose that Assumption 8.1 holds and let $\mathcal{M}$ defines the NCM of Definition 6.2, which models $M$ of the CV-STEM contraction metric in Theorem 4.2 for the nominal system (8.1) with $\theta=\theta_{n}$, constructed with an additional convex constraint given as $\partial_{b_{i}(x)}\bar{W}+\partial_{b_{i}(x_{d})}\bar{W}=0$, where $\partial_{b_{i}(q)}\bar{W}=\sum_{i}(\partial\bar{W}/\partial q_{i})b_{i}(q)$ for $B=[b_{1},\cdots,b_{m}]$ (see Manchester and Slotine [2017], Lopez and Slotine [2021]). Suppose also that the matched uncertainty condition Lopez and Slotine [2021] holds, i.e., $(\Pi(x)-\Pi(x_{d}))^{\top}\vartheta\in\textrm{span}(B(x))$ for $\Pi(x)$, and that (8.1) is controlled by the following adaptive control law:
@@
where $u_{L}$ is given by (6.7) of Theorem 6.3, i.e., $u_{L}=u_{d}-R^{-1}B^{\top}\mathcal{M}(x-x_{d})$, $\Gamma\in\mathbb{R}^{{c}\times{c}}$ is a diagonal matrix with positive elements that governs the rate of adaptation, $\sigma\in\mathbb{R}_{\geq 0}$, and $(\Pi(x)-\Pi(x_{d}))^{\top}\vartheta=B(x)\varphi(x,x_{d})^{\top}\vartheta$.
If $\exists\underline{\gamma},\overline{\gamma},\bar{b},\bar{\rho},\bar{\phi},\bar{\vartheta}\in\mathbb{R}_{>0}$ s.t. $\underline{\gamma}\mathrm{I}\preceq\Gamma\preceq\overline{\gamma}\mathrm{I}$, $\|B(x)\|\leq\bar{b}$, $\|R^{-1}(x,x_{d},u_{d})\|\leq\bar{\rho}$, $\|\varphi(x,x_{d})\|\leq\bar{\phi}$, and $\|\vartheta\|\leq\bar{\vartheta}$, and if $\Gamma$ and $\sigma$ of (8.4) are selected to satisfy the following relation for the learning error $\|\mathcal{M}-M\|\leq\epsilon_{\ell}$ in some compact set $\mathcal{S}$ as in (6.22) of Theorem 6.3:
@@
for $\exists\alpha_{a}\in\mathbb{R}_{>0}$, $\alpha_{\ell}$ given in Theorem 6.3, and $\underline{m}$ and $\overline{m}$ given in $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), then the system (8.1) is robust against bounded deterministic and stochastic disturbances with $\sigma\neq 0$, and we have the following bound in the compact set $\mathcal{S}$:
@@
where $\mathtt{e}=x-x_{d}$, and $V_{\ell}=\int^{\xi_{1}}_{\xi_{0}}\|\Theta\delta q\|$ is defined in Theorem 2.3 with $M=\Theta^{\top}\Theta$ replaced by $\mathrm{diag}(M,\Gamma^{-1})$ for $\xi_{0}=[x_{d}^{\top},\vartheta^{\top}]^{\top}$ and $\xi_{1}=[x^{\top},\hat{\vartheta}^{\top}]^{\top}$. Furthermore, if the learning error $\epsilon_{\ell}=0$ (CV-STEM control), (8.4) with $\sigma=0$ guarantees asymptotic stability of $x$ to $x_{d}$ in (8.1).
@@
The proof can be found in Tsukamoto et al. [2021], but here we emphasize the use of contraction theory. For $u$ given by (8.3), the virtual system of a smooth path $q(\mu,t)=[q_{x}^{\top},q_{\vartheta}^{\top}]^{\top}$ parameterized by $\mu\in[0,1]$, which has $q(\mu=0,t)=[x_{d}^{\top},\vartheta^{\top}]^{\top}$ and $q(\mu=1,t)=[x^{\top},\hat{\vartheta}^{\top}]^{\top}$ as its particular solutions, is given as follows:
@@
The arguments of $A(\varrho,x,x_{d},u_{d})$, $B(x)$, $\mathcal{M}(x,x_{d},u_{d})$, and $\varphi(x,x_{d})$ are omitted for notational simplicity. Since $\dot{q}_{x}=\zeta(q_{x},x,x_{d},u_{d},t)$ is contracting due to Theorems 4.2 and 6.3 with a contraction rate given by $\alpha_{\ell}$ in Theorem 6.3, we have for a Lyapunov function $V=\delta q_{x}^{\top}M\delta q_{x}+\delta q_{\vartheta}^{\top}\Gamma^{-1}\delta q_{\vartheta}$ that
@@
Since we have $\|\partial d_{q\vartheta}/\partial\mu\|=\sigma\bar{\vartheta}$, this implies $\dot{V}_{\ell}\leq-\alpha_{a}V_{\ell}+\sigma\sqrt{\overline{\gamma}}\bar{\vartheta}$, yielding the bound (8.6) due to Lemma 2.1 Tsukamoto et al. [2021]. Robustness against deterministic and stochastic disturbances follows from Theorem 3.1 if $\sigma\neq 0$. If $\epsilon_{\ell}=0$ and $\sigma=0$, the relation (8.9) reduces to $\dot{V}/2\leq-\alpha\delta q_{x}^{\top}M\delta q_{x}$, which results in asymptotic stability of $x$ to $x_{d}$ in (8.1) due to Barbalat’s lemma [Khalil, 2002, p. 323] as in the proof of Theorem 2 in Lopez and Slotine [2021]. $\qed$
@@
Although Theorem 8.1 is for the case where $f(x)$ is affine in its parameter, it is also useful for the following types of systems with an uncertain parameter $\theta\in\mathbb{R}^{c}$ and a control input $u$ (see Tsukamoto et al. [2021]):
@@
where ${p}\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{n}$, $h:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $H:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$, $\Pi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times{c}}$, $x=[({p}^{(n-2)})^{\top},\cdots,({p})^{\top}]^{\top}$, and ${p}^{(k)}$ denotes the $k$th time derivative of p{p}. In particular, adaptive sliding control Slotine and Coetsee [1986] designs $u$ to render the system of the composite variable $s$ given as $s={p}^{(n-1)}-{p}_{r}^{(n-1)}$ to be contracting, where ${p}_{r}^{(n-1)}=p_{d}^{(n-1)}-\sum_{i=0}^{n-2}\lambda_{i}\mathtt{e}^{(i)}$, $\mathtt{e}=p-p_{d}$, $p_{d}$ is a target trajectory, and $\kappa^{n-1}+\lambda_{n-2}\kappa^{n-2}+\cdots+\lambda_{0}$ is a stable (Hurwitz) polynomial in the Laplace variable $\kappa$ (see Example 2.6). Since we have $\mathtt{e}^{(n-1)}=s-\sum_{i=0}^{n-2}\lambda_{i}\mathtt{e}^{(i)}$ and the system for $\mathtt{e}^{(0)},\cdots,\mathtt{e}^{(n-2)}]$ is also contracting if $s=0$ due to the Hurwitz property, the hierarchical combination property Slotine [2003], Slotine and Lohmiller [2001] of contraction in Theorem 2.7 guarantees $\lim_{t\to\infty}\|p-p_{d}\|=0$ [Slotine and Li, 1991, p. 352] (see Example 2.7).
@@
Using the technique of Remark 8.1, we can construct adaptive control for the Lagrangian system in Example 2.6 as follows:
@@
where $\hat{\mathcal{H}}$, $\hat{\mathcal{C}}$, and $\hat{\mathcal{G}}$ are the estimates of $\mathcal{H}$, $\mathcal{C}$, and $\mathcal{G}$, respectively, and the other variables are as defined in Example 2.6. Suppose that the terms $\mathcal{H}$, $\mathcal{C}$, and $\mathcal{G}$ depend linearly on the unknown parameter vector $\theta$ as follows [Slotine and Li, 1991, p. 405]:
@@
which results in asymptotic stability of the differential state $\delta q$ (i.e., semi-contraction Wang and Slotine [2005], see also Barbalat’s lemma [Slotine and Li, 1991, p. 405-406]).
8.2 Parameter-Dependent Contraction Metric (aNCM)
Although Theorem 8.1 utilizes the NCM designed for the nominal system (8.1) with $\theta=\theta_{n}$, we could improve its representational power by explicitly taking the parameter estimate $\hat{\theta}$ as one of the NCM arguments Tsukamoto et al. [2021], leading to the concept of an adaptive NCM (aNCM).
In this section, we consider multiplicatively-separable nonlinear systems given in Assumption 8.2, which holds for many types of systems including robotics systems Slotine and Li [1991], spacecraft high-fidelity dynamics Battin [1999], Morgan et al. [2012], and systems modeled by basis function approximation and DNNs Nelles [2001], Sanner and Slotine [1992].

### v8

@@ 7.3. Tube-based State Constraint Satisfaction
7.4. Safe Exploration and Learning of Disturbances
All the theorems so far assume that the sizes of unknown external disturbances are fixed (i.e., $\bar{d}$ in Theorem LABEL:THM:Thm:Robust_contraction_original and $(\bar{g}_{0},\bar{g}_{1})$ in Theorem LABEL:THM:Thm:robuststochastic are known and fixed). Since such an assumption could yield conservative state constraints, e.g., in utilizing motion planning of Lemma 7.1 with the tube of Theorem LABEL:THM:Thm:lagros_stability, we could consider better estimating the unknown parts, $d(x,t)$ in Theorem LABEL:THM:Thm:Robust_contraction_original or $G_{0}(x,t)$ and $G_{1}(x,t)$ in Theorem LABEL:THM:Thm:robuststochastic, also using contraction theory and machine learning. For example, it is demonstrated in 9290355 that the stochastic bounds of Theorem LABEL:THM:Thm:robuststochastic can be used to ensure safe exploration and reduction of uncertainty over epochs in learning unknown, control non-affine residual dynamics. Here, its optimal motion plans are designed by solving chance-constrained trajectory optimization and executed using a feedback controller such as Theorem LABEL:THM:Thm:CV-STEM with a control barrier function-based safety filter 7040372, 8405547. It is also shown in 9303957, L1contraction2, sun2021uncertaintyaware that the state-dependent disturbance $d(x)$ in Theorem LABEL:THM:Thm:robuststochastic can be learned adaptively with the Gaussian process while ensuring safety.
These techniques are based on the robustness and stability properties of contraction theory, which guarantee their state trajectories to stay in a tube around the target trajectory as in Theorems LABEL:THM:Thm:Robust_contraction_originalLABEL:THM:Thm:robuststochasticLABEL:THM:Thm:contraction_learning, and LABEL:THM:Thm:contraction_learning_sto, and at the same time, enable safely sampling training data for learning the unknown part of the dynamics. Sections 8 and 9 are for giving an overview of these concepts in the context of contraction theory-based adaptive control ancm, lopez2021universal, 9109296 and robust control synthesis for learned models neurallander, 47710, boffi2020learning, cdc_systemid, 7992901.
@@ 7.4. Safe Exploration and Learning of Disturbances
8. Learning-based Adaptive Control
@@
where $x$, $u$, $x_{d}$, and $u_{d}$ are as defined in (6.1) and (6.3), and ${f}:\mathbb{R}^{n}\times\mathbb{R}^{c}\mapsto\mathbb{R}^{n}$ and $B:\mathbb{R}^{n}\times\mathbb{R}^{c}\mapsto\mathbb{R}^{n\times m}$ are known smooth functions with the uncertain parameter $\theta\in\mathbb{R}^{c}$. Due to Theorems LABEL:THM:Thm:Robust_contraction_original and LABEL:THM:Thm:contraction_learning, we can see that the robust control techniques presented earlier in Theorems LABEL:THM:Thm:CV-STEMLABEL:THM:Thm:ccm_cvstemLABEL:THM:Thm:ncm_clf, and LABEL:THM:Thm:ncm_ccm_clf are still useful in this case if the modeling errors $\|f(x,\theta)-f(x,\theta_{n})\|$ and $\|B(x,\theta)-B(x,\theta_{n})\|$ are bounded, where $\theta_{n}$ is a nominal guess of the true parameter $\theta$. However, there are situations where such assumptions are not necessarily true.
We present a method of deep learning-based adaptive control for nonlinear systems with parametric uncertainty, thereby further improving the real-time performance of robust control in Theorems LABEL:THM:Thm:contraction_learning and LABEL:THM:Thm:contraction_learning_sto for model-based systems, and Theorems LABEL:THM:Thm:NCMstability_modelfree and LABEL:THM:Thm:neurallander for model-free systems. Although we consider continuous-time dynamical systems in this section, discrete-time changes could be incorporated in this framework using regretboffi, 1469901. Also, note that the techniques in this section ancm can be used with differential state feedback frameworks ccm, 7989693, 47710, WANG201944, vccm, rccm as described in Theorem LABEL:THM:Thm:ccm_cvstem, trading off added computational cost for generality (see Table 3 and 9109296, lopez2021universal).
8.1. Adaptive Control with CV-STEM and NCM
@@
leading to the following theorem ancm for the NCM-based adaptive control. Note that we could also use the SDC formulation with respect to a fixed point as delineated in Theorem LABEL:THM:Thm:fixed_sdc mypaperTAC, Ref:Stochastic, mypaper.
@@
Learning-based Adaptive Control with Contraction TheoryThm:adaptive_robust_affine Suppose that Assumption 8.1 holds and let $\mathcal{M}$ defines the NCM of Definition LABEL:DEF:Def:NCM, which models $M$ of the CV-STEM contraction metric in Theorem LABEL:THM:Thm:CV-STEM for the nominal system (8.1) with $\theta=\theta_{n}$, constructed with an additional convex constraint given as $\partial_{b_{i}(x)}\bar{W}+\partial_{b_{i}(x_{d})}\bar{W}=0$, where $\partial_{b_{i}(q)}\bar{W}=\sum_{i}(\partial\bar{W}/\partial q_{i})b_{i}(q)$ for $B=[b_{1},\cdots,b_{m}]$ (see ccm, 9109296). Suppose also that the matched uncertainty condition 9109296 holds, i.e., $(\Pi(x)-\Pi(x_{d}))^{\top}\vartheta\in\textrm{span}(B(x))$ for $\Pi(x)$, and that (8.1) is controlled by the following adaptive control law:
@@
where $u_{L}$ is given by (6.7) of Theorem LABEL:THM:Thm:NCMstability2, i.e., $u_{L}=u_{d}-R^{-1}B^{\top}\mathcal{M}(x-x_{d})$, $\Gamma\in\mathbb{R}^{{c}\times{c}}$ is a diagonal matrix with positive elements that governs the rate of adaptation, $\sigma\in\mathbb{R}_{\geq 0}$, and $(\Pi(x)-\Pi(x_{d}))^{\top}\vartheta=B(x)\varphi(x,x_{d})^{\top}\vartheta$.
If $\exists\underline{\gamma},\overline{\gamma},\bar{b},\bar{\rho},\bar{\phi},\bar{\vartheta}\in\mathbb{R}_{>0}$ s.t. $\underline{\gamma}\mathrm{I}\preceq\Gamma\preceq\overline{\gamma}\mathrm{I}$, $\|B(x)\|\leq\bar{b}$, $\|R^{-1}(x,x_{d},u_{d})\|\leq\bar{\rho}$, $\|\varphi(x,x_{d})\|\leq\bar{\phi}$, and $\|\vartheta\|\leq\bar{\vartheta}$, and if $\Gamma$ and $\sigma$ of (8.4) are selected to satisfy the following relation for the learning error $\|\mathcal{M}-M\|\leq\epsilon_{\ell}$ in some compact set $\mathcal{S}$ as in (6.22) of Theorem LABEL:THM:Thm:NCMstability2:
@@
for $\exists\alpha_{a}\in\mathbb{R}_{>0}$, $\alpha_{\ell}$ given in Theorem LABEL:THM:Thm:NCMstability2, and $\underline{m}$ and $\overline{m}$ given in $\underline{m}\mathrm{I}\preceq M\preceq\overline{m}\mathrm{I}$ of (2.32), then the system (8.1) is robust against bounded deterministic and stochastic disturbances with $\sigma\neq 0$, and we have the following bound in the compact set $\mathcal{S}$:
@@
where $\mathtt{e}=x-x_{d}$, and $V_{\ell}=\int^{\xi_{1}}_{\xi_{0}}\|\Theta\delta q\|$ is defined in Theorem LABEL:THM:Thm:path_integral with $M=\Theta^{\top}\Theta$ replaced by $\mathrm{diag}(M,\Gamma^{-1})$ for $\xi_{0}=[x_{d}^{\top},\vartheta^{\top}]^{\top}$ and $\xi_{1}=[x^{\top},\hat{\vartheta}^{\top}]^{\top}$. Furthermore, if the learning error $\epsilon_{\ell}=0$ (CV-STEM control), (8.4) with $\sigma=0$ guarantees asymptotic stability of $x$ to $x_{d}$ in (8.1).
@@
The proof can be found in ancm, but here we emphasize the use of contraction theory. For $u$ given by (8.3), the virtual system of a smooth path $q(\mu,t)=[q_{x}^{\top},q_{\vartheta}^{\top}]^{\top}$ parameterized by $\mu\in[0,1]$, which has $q(\mu=0,t)=[x_{d}^{\top},\vartheta^{\top}]^{\top}$ and $q(\mu=1,t)=[x^{\top},\hat{\vartheta}^{\top}]^{\top}$ as its particular solutions, is given as follows:
@@
The arguments of $A(\varrho,x,x_{d},u_{d})$, $B(x)$, $\mathcal{M}(x,x_{d},u_{d})$, and $\varphi(x,x_{d})$ are omitted for notational simplicity. Since $\dot{q}_{x}=\zeta(q_{x},x,x_{d},u_{d},t)$ is contracting due to Theorems LABEL:THM:Thm:CV-STEM and LABEL:THM:Thm:NCMstability2 with a contraction rate given by $\alpha_{\ell}$ in Theorem LABEL:THM:Thm:NCMstability2, we have for a Lyapunov function $V=\delta q_{x}^{\top}M\delta q_{x}+\delta q_{\vartheta}^{\top}\Gamma^{-1}\delta q_{\vartheta}$ that
@@
Since we have $\|\partial d_{q\vartheta}/\partial\mu\|=\sigma\bar{\vartheta}$, this implies $\dot{V}_{\ell}\leq-\alpha_{a}V_{\ell}+\sigma\sqrt{\overline{\gamma}}\bar{\vartheta}$, yielding the bound (8.6) due to Lemma 2.1 ancm. Robustness against deterministic and stochastic disturbances follows from Theorem LABEL:THM:Thm:CVSTEM:LMI if $\sigma\neq 0$. If $\epsilon_{\ell}=0$ and $\sigma=0$, the relation (8.9) reduces to $\dot{V}/2\leq-\alpha\delta q_{x}^{\top}M\delta q_{x}$, which results in asymptotic stability of $x$ to $x_{d}$ in (8.1) due to Barbalat’s lemma [Khalil:1173048, p. 323] as in the proof of Theorem 2 in 9109296. $\qed$
@@
Although Theorem LABEL:THM:Thm:adaptive_robust_affine is for the case where $f(x)$ is affine in its parameter, it is also useful for the following types of systems with an uncertain parameter $\theta\in\mathbb{R}^{c}$ and a control input $u$ (see ancm):
@@
where ${p}\in\mathbb{R}^{n}$, $u\in\mathbb{R}^{n}$, $h:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$, $H:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times n}$, $\Pi:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times{c}}$, $x=[({p}^{(n-2)})^{\top},\cdots,({p})^{\top}]^{\top}$, and ${p}^{(k)}$ denotes the $k$th time derivative of p{p}. In particular, adaptive sliding control adaptive_sliding designs $u$ to render the system of the composite variable $s$ given as $s={p}^{(n-1)}-{p}_{r}^{(n-1)}$ to be contracting, where ${p}_{r}^{(n-1)}=p_{d}^{(n-1)}-\sum_{i=0}^{n-2}\lambda_{i}\mathtt{e}^{(i)}$, $\mathtt{e}=p-p_{d}$, $p_{d}$ is a target trajectory, and $\kappa^{n-1}+\lambda_{n-2}\kappa^{n-2}+\cdots+\lambda_{0}$ is a stable (Hurwitz) polynomial in the Laplace variable $\kappa$ (see Example LABEL:EX:ex:lag_metric). Since we have $\mathtt{e}^{(n-1)}=s-\sum_{i=0}^{n-2}\lambda_{i}\mathtt{e}^{(i)}$ and the system for $\mathtt{e}^{(0)},\cdots,\mathtt{e}^{(n-2)}]$ is also contracting if $s=0$ due to the Hurwitz property, the hierarchical combination property Ref:contraction2, Ref:contraction5 of contraction in Theorem LABEL:THM:Thm:Robust_contraction_hierc guarantees $\lim_{t\to\infty}\|p-p_{d}\|=0$ [Ref_Slotine, p. 352] (see Example LABEL:EX:ex:hiera).
@@
Adaptive Control of Lagrangian Systemsex:lag_adaptive0 Using the technique of Remark 8.1, we can construct adaptive control for the Lagrangian system in Example LABEL:EX:ex:lag_metric as follows:
@@
where $\hat{\mathcal{H}}$, $\hat{\mathcal{C}}$, and $\hat{\mathcal{G}}$ are the estimates of $\mathcal{H}$, $\mathcal{C}$, and $\mathcal{G}$, respectively, and the other variables are as defined in Example LABEL:EX:ex:lag_metric. Suppose that the terms $\mathcal{H}$, $\mathcal{C}$, and $\mathcal{G}$ depend linearly on the unknown parameter vector $\theta$ as follows [Ref_Slotine, p. 405]:
@@
which results in asymptotic stability of the differential state $\delta q$ (i.e., semi-contraction Ref:contraction3, see also Barbalat’s lemma [Ref_Slotine, p. 405-406]).
8.2. Parameter-Dependent Contraction Metric (aNCM)
Although Theorem LABEL:THM:Thm:adaptive_robust_affine utilizes the NCM designed for the nominal system (8.1) with $\theta=\theta_{n}$, we could improve its representational power by explicitly taking the parameter estimate $\hat{\theta}$ as one of the NCM arguments ancm, leading to the concept of an adaptive NCM (aNCM).
In this section, we consider multiplicatively-separable nonlinear systems given in Assumption 8.2, which holds for many types of systems including robotics systems Ref_Slotine, spacecraft high-fidelity dynamics battin, doi:10.2514/1.55705, and systems modeled by basis function approximation and DNNs Nelles2001, SannerSlotine1992.